OpenStudy (anonymous):

will fan & medal!!! PLEASE HELP ME! Determine which ordered pair is a solution of y=x^2-2. a. (-2, 6) b. (5, 21) c. (-1, -1) d. (-5, 24)

OpenStudy (anonymous):

ok so what do you think it is

rishavraj (rishavraj):

just plug the values given in the options and see whether it satisfies the equation or not

OpenStudy (anonymous):

this is acully not the best way

rishavraj (rishavraj):

hmmmm see ....like option A is (-2, 6) ....means plug x = -2 and y = 6 and this doesn't satisfy the equation......likewise check other options

OpenStudy (anonymous):

ohhhh...let me see

rishavraj (rishavraj):

take ur time :)) @-minibubbles-

OpenStudy (butterflydreamer):

i believe the best way to solve this question is to do what @rishavraj said :) Simply plug in the values from your options and see which one works! (LHS = RHS) So for the first one.. point (-2 ,6 ) Your x = -2 , y = 6 Sub these values into your equation " y=x^2-2" Basically replace x with -2 and y with 6 and we get: 6 = (-2)^2 - 2 6 = 4 - 2 6 = 2 which is incorrect so not this one :) It doesn't satisfy the equation.